OpenStudy (anonymous):

Christina bought a yoyo from a company that claims that, with each retraction, the string rolls up by 70% of the original length. She sets up a tape measure and throws the yoyo 3 times. Her data are charted below. Throw Length of string (feet) 1 3 2 2.1 3 1.47 Christina wants to find the sum of the length of string after 10 throws. What is the sum of the lengths, rounded to the nearest hundredth?

OpenStudy (tkhunny):

1) Does70% prove to be correct? 2) Break out your best Geometric Series sum formula.

OpenStudy (anonymous):

idk??

OpenStudy (tkhunny):

Well, why not? What's 70% of 3?

OpenStudy (anonymous):

2.1

OpenStudy (anonymous):

ahhh 0.121060821

OpenStudy (tkhunny):

Okay, you now have 10 throws to throw. 3 + 3*.7 + 3*.7^2 +...+ 3*.7^9 Is that 10 throws? You have to believe.

OpenStudy (anonymous):

I believe ...!! praise the LAWD

OpenStudy (anonymous):

OMG THE SUM

OpenStudy (tkhunny):

Add them up: \(\dfrac{3 - 3*0.7^{10}}{1-0.7}\) This should be looking somewhat familiar.

OpenStudy (anonymous):

YES, IT DOES

OpenStudy (ckaranja):

This is a geometric series that follows a decreasing order. He simply should use \[ar^{n-1}\] In this case a=3, r=\[\frac{ 2.1 }{ 3 }\] and n=10 So, \[(3*\frac{ 2.1 }{ 3})^{9}\]

OpenStudy (tkhunny):

@ckaranja Except that you probably should reread the problem statement.